OpenStudy (abhisar):

Can some body make me understand this? \[\frac{d T}{d (1/T)} = \frac{1}{\left(\frac{d (1/T)}{d T}\right)} \] How?

OpenStudy (abhisar):

@ganeshie8 @myininaya

Elsa213 (elsa213):

@Hero

myininaya (myininaya):

\[\frac{dT}{d(\frac{1}{T})} \\ \text{ \let } u=\frac{1}{T} \text{ then } T=\frac{1}{u} \\ \text{ so } \frac{dT}{d(\frac{1}{T})} =\frac{d(\frac{1}{u})}{du} \cdot \frac{du}{dT}\]

myininaya (myininaya):

\[\frac{1}{\frac{d(\frac{1}{T})}{dT}}=\frac{dT}{d(\frac{1}{T})} \text{ since the reciprocal of } \frac{d(\frac{1}{T})}{dT} \text{ is } \frac{dT}{d(\frac{1}{T})}\]

myininaya (myininaya):

wait what do you want to understand about this?

OpenStudy (abhisar):

Like how can you write \[\frac{d T}{d (1/T)}~as~\frac{1}{\left(\frac{d (1/T)}{d T}\right)} \]

OpenStudy (abhisar):

Have you used chain rule or something?

myininaya (myininaya):

yes I used chain rule above

OpenStudy (abhisar):

Umm..But I thought chain rule is used when we differentiate a function with respect to different function. Right?

myininaya (myininaya):

well I did introduce a new function

myininaya (myininaya):

would you agree that: \[\frac{a}{b}=\frac{1}{\frac{b}{a}}?\]

OpenStudy (abhisar):

Yes.

OpenStudy (abhisar):

Whoops!!!!!!! lol, I was thinking it to be some complex stuff.

OpenStudy (abhisar):

Thanks a bunch c:

myininaya (myininaya):

np